When making a telephone call using a calling card, a call lasting 6 minutes costs $2.90. A call lasting 16 minutes cost $6.90. Find a linear function that relates the cost, C, of a making a call length, X, of the call, treating length of the call as the independent variable.
Answer by stanbon(75887)

When making a telephone call using a calling card, a call lasting 6 minutes costs $2.90. A call lasting 16 minutes cost $6.90. Find a linear function that relates the cost, C, of a making a call length, X, of the call, treating length of the call as the independent variable.
---------------------------
You have two points : (6,2.9) and (16,6.9)
-----
slope = (6.9-2.9)/(16-6) = 4/10 = 2/5
-----
Linear Form: cost = (slope)(length of call)+b
Solve for "b":
2.9 = (2/5)(6)+b
b = 2.9-(2/5)(6)
b = 0.5
-----------------------
Equation:
cost = (2/5)x + (1/2)
===========================
Cheers,
Stan H.